Operations Report Nov 27th – Dec 03rd

MDRS Operations Report 30-NOV-2021- 3 Dec-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Charge controller problem resolved

Action taken for non-nominal systems: I was able to reset on the second try

Generator: The electrician who wired the system together originally will be coming to figure out the problem on either the 13th or 14th. We had a rather intimidating experience with the system and I didn’t want us to get hurt with my experiments. I can’t fix this one.

ScienceDome Dual Split: Off

Solar—Nominal, supplying all power

Solar— VDC Last 24 hours:

Average 50.67

Minimum 45.90

Maximum 58.40

Notes on solar: These numbers are reasonable with the addition of the heat in the Hab and lights in the GreenHab,

Propane Reading, station tank – FULL %

Propane Reading, director tank— FULL %

Propane Reading, intern tank— 50 %

Propane Reading, generator— 80 %

Ethanol Free Gasoline – 1 gallons

Water (loft tank): 40 gallons

Water Meter: not checked units

Water (static tank) – 300

Static to Loft Pump used – yes

Water in GreenHab – 100 gallons (not accessible)

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 150 gallons

Hab toilet tank emptied: no

Perseverance rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: No

Sojourner rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: Yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: No

Opportunity rover used: Yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: No

Curiosity rover used: Yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: We are unplugging the rovers at night on a rotating basis and only charging them in the daytime

ATV’s Used: (Honda, 350.1, 350.2, 300): Honda’s new tires came but one of them went flat and was returned and fixed. We should get them on and the machine operational tomorrow.

Reason for use: n/a

Oil Added? No

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town to get water and run errands.

CrewCar used and why, where? Yes, to Green River to return Allan to the station and to Bicknell for me to get my COVID booster

Luna used and why, where? Yes, to pick up Atila, Allan and Sebastian in Green River

General notes and comments: CrewCar needs a new battery before the crews return but is fine when run everyday. The tow lights have disconnected from the CrewCar hitch and will be repaired tomorrow.

Summary of internet: All accounts are nominal

Summary of suits and radios: All unplugged except for communication radios

Campus wide inspection, if action taken, what and why? : Nothing to report

Summary of general operations: All propane tanks have been switched and we are awaiting pick up of the old tanks, which are currently stored next to the Hab tank.

Summary of Hab operations: Atila and Sebastian are crewing the Hab while we finish the intern trailer. We have two more interns arriving this weekend and they will stay there too.

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: Plants were fertilized. Supplemental light with old fixtures from 7 pm-midnight.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Nothing at this time.